# Unified Taxonomy for Multivariate Time Series Anomaly Detection using Deep Learning

### 19 Mar 2026 v1

@@ 5.3 Future research directions and open challenges
Future MTSAD research is moving toward adaptive, interpretable and scalable spatio-temporal architectures. Graph-Transformer hybrids such as LGAT exemplify the value of unifying attention-based temporal modeling with dynamic graph reasoning for non-stationary data [LGAT]. Generative models like MAFCD (Multi-level and Adaptive Conditional Diffusion) emphasize multi-view probabilistic learning for robust anomaly synthesis [MAFCD], while memory-augmented designs such as MemMambaAD enhance long-term contextual memory [MemMambaAD]. Additionally, explainable models like SHAPAttenVAE (Shapley Attention VAE) signal growing attention to interpretability and causal reasoning in the field [SHAPAttenVAE].
@@
Overall, this study confirms that DL-based methods currently dominate the MTSAD landscape. However, some authors question whether this paradigm represents the optimal path forward [Audibert2022, Wu2021]. Although DL methods tend to handle large datasets more efficiently than conventional or ML-based approaches, there is no conclusive evidence that the first consistently outperform the latter across a wide range of datasets and real-world conditions. Moreover, reproducibility remains a major obstacle for adoption in real scenarios, as multiple factors, including model design, hyperparameter settings and computational resources influence results. These challenges suggest that the field should not rely exclusively on DL-based paradigms. In certain applications, classical approaches, ML-based methods or hybrid frameworks may offer more suitable and interpretable solutions. Furthermore, integrating domain knowledge through expert systems into data-driven models represents a promising direction, combining the strengths of artificial intelligence with human expertise to improve robustness, reliability and applicability in real-world anomaly detection tasks.
@@
One final note concerns the evaluation of MTSAD methods, which relies heavily on a limited set of benchmark datasets. This review does not detail benchmark datasets, as this has been covered elsewhere [ZamanzadehDarban2024, Garg2022]. Nonetheless, concerns remain about the reliability of widely used datasets such as MSL [MSL], SMAP [MSL] and SMD [OmniAnomaly]. Wu2021 argue that these benchmarks may create an illusion of progress due to trivial anomalies, mislabeled instances and unrealistic conditions. The field thus requires new, high-quality and diverse datasets to enable a fair comparison between the approaches and a genuine quantification of the advancements of the research field.

### 24 Apr 2026 v2

@@ 5.3 Future research directions and open challenges
Building on the analysis presented throughout this review, several open challenges and directions emerge that are likely to shape the future of MTSAD research. In particular, three main directions can be identified
  • Adaptive, interpretable and scalable spatio-temporal architectures;
  • Moving beyond exclusively DL-based approaches for MTSAD;
  • Awareness of the limitations of existing benchmarks.
Each of these directions is discussed below.
@@
The first direction concerns the development of adaptive, interpretable and scalable spatio-temporal architectures for MTSAD. Graph-Transformer hybrids such as LGAT exemplify the value of unifying attention-based temporal modeling with dynamic graph reasoning for non-stationary data [LGAT]. Generative models like MAFCD (Multi-level and Adaptive Conditional Diffusion) emphasize multi-view probabilistic learning for robust anomaly synthesis [MAFCD], while memory-augmented designs such as MemMambaAD enhance long-term contextual memory [MemMambaAD]. Additionally, explainable models like SHAPAttenVAE (Shapley Attention VAE) signal growing attention to interpretability and causal reasoning in the field [SHAPAttenVAE]. Real-world deployment demands models that go beyond accuracy, requiring transparency, adaptability to non-stationary environments and computational efficiency. Therefore, these directions are likely to become central to the next generation of MTSAD research.
@@
The second direction reflects a growing recognition of the need to move beyond exclusively DL-based approaches for MTSAD. While DL-based methods currently dominate the field, several studies question whether this paradigm consistently represents the most effective solution across diverse settings [Audibert2022, Wu2021]. Although DL methods tend to handle large datasets more efficiently than conventional or ML-based approaches, there is no conclusive evidence that the first consistently outperform the latter across a wide range of datasets and real-world conditions. Moreover, reproducibility remains a major obstacle for adoption in real scenarios, as multiple factors, including model design, hyperparameter settings and computational resources influence results. These challenges suggest that the field should not rely exclusively on DL-based paradigms. In certain applications, classical approaches, ML-based methods or hybrid frameworks may offer more suitable and interpretable solutions. Furthermore, integrating domain knowledge through expert systems into data-driven models represents a promising direction, combining the strengths of artificial intelligence with human expertise to improve robustness, reliability and applicability in real-world anomaly detection tasks. Rethinking the DL-only paradigm is critical, as increasing model complexity does not guarantee better performance in practice and may further distance research from real-world applicability.
@@
The third direction relates to the limitations of current benchmark datasets and evaluation protocols. Although this review does not provide a detailed analysis of benchmark datasets, as this topic has been extensively covered in prior work [ZamanzadehDarban2024, Garg2022], it is important to acknowledge ongoing concerns regarding their reliability and representativeness. Nonetheless, concerns remain about the reliability of widely used datasets such as MSL [MSL], SMAP [MSL] and SMD [OmniAnomaly]. Wu2021 argue that these benchmarks may create an illusion of progress due to trivial anomalies, mislabeled instances and unrealistic conditions. Without reliable and diverse benchmarks, it becomes difficult to determine whether improvements reflect genuine advances or simply overfitting to the particularities of widely used datasets. Therefore, the development of new, high-quality and diverse datasets is essential to enable fair comparisons between approaches and genuine quantification of progress in the field.